A beaker with 1.90x102 mL of an acetic acid buffer with a pH of 5.000 is sitting on a benchtop. The total 0.460 MHCl solution to the beaker. How much will the pH change? The pKa of acetic acid is 4.740 Express your answer numerically to two decimal places. Use a minus (-sign if the pH has decreased.
Answer #1

pH of acidic buffer = pka + log(acetate/acetic acid)

pka of acetic acid = 4.74

Total no of mol of buffer = 0.1*190 = 19 mmol

acetate + acetic acid = 19 mmol


5 = 4.74+log(x/(19-x))

x = acetate = 12 mmol

   aceticacid = 19-12 = 7 mmol

No of mol of HCl added = 7.6*0.46 = 3.5 mmol

pH of acidic buffer = pka + log(acetate-HCl/acetic acid+HCl)

                    = 4.74+log((12-3.5)/(7+3.5))

                    = 4.65
so that, pH decreases

change in pH = 4.65-5 = -0.35